OpenStudy (anonymous):

What is the radius of the circle whose equation is x2 + y2 = 8?

OpenStudy (anonymous):

\[2\sqrt{2}\] 4 64

OpenStudy (phi):

do you know the equation for a circle ? (look in your notes) what is it ?

OpenStudy (igreen):

Easiest way is to graph.. https://www.desmos.com/calculator Then see where the farthest point the circle goes to the left/right.

OpenStudy (anonymous):

\[x ^{2}+y^{2}=8\]

OpenStudy (phi):

I mean what is the equation for any circle ? see http://www.mathwarehouse.com/geometry/circle/equation-of-a-circle.php

OpenStudy (phi):

you should match up your equation with \[ x^2 +y^2 = r^2 \] and see that \[ r^2 = 8 \] to find r, take the square root \[ r = \sqrt{8} \]

OpenStudy (phi):

you can "simplify" the square root of 8 by writing 8 as 2*2*2 \[ \sqrt{2 \cdot 2 \cdot 2} = ? \]

OpenStudy (anonymous):

2*2*2=8

OpenStudy (phi):

yes sqr(8) can be written as sqr(2*2*2)= sqr(2*2) * sqr(2)

OpenStudy (igreen):

Right. so going with what @phi just said, you will end up with \(2 \sqrt 2\).